OpenStudy (anonymous):

The sum of 2 numbers is 18. The difference between 3 times the smaller number and the larger number is 10. Define variables and formulate a system of equations from this situation. What is the larger number? work please? :D

OpenStudy (anonymous):

x+y=18 3x-y=10 add both equations to get 4x=28 x=7 y=18-x= 18-7=11 y=11